OpenStudy (anonymous):

The pole for a tent makes a 59 degree angle with the ground. If the load on the pole is 15 pounds, what are the vertical and horizontal loads?

OpenStudy (anonymous):

Draw a right triangle like an "uphill" if you were walking from left to right. Label the long side 15 lbs. Label the vertical side y. Label the horizontal side x. Using SOH CAH TOA, sin 59 = opposite / hypotenuse = y / 15 y = 15 sin 59 = 12.858 lbs cos 59 = adjacent / hypotenuse = x / 15 x = 15 cos 59 = 7.726 lbs Answer: Horizontal Load: 7.73 lbs Vertical Load: 12.86 lbs